OpenStudy (anonymous):

solve for x: log2 (x-3) + log2 (x+1) = 5

OpenStudy (anonymous):

I suggest using wolfram alpha, just search it in your browser

OpenStudy (anonymous):

thank you!

OpenStudy (el_tucan):

this is the property: logb(xy) = logbx + logby you multiply (x-3) & (x+1)

OpenStudy (anonymous):

start with \[\log_2((x-3)(x+1))=5\] and then rewrite in exponential form as \[(x-3)(x+1)=2^5\] solve the quadratic, check your solutions